$ 1 yh

What is the total Rt (ohms), current I, V1, V2, and V3?
Answer #1

Ans.)

Using krichhoff voltage law in given circuit we get:-

16-12 = 5i + i + 2i + 4i

4 = 12i

I = (1/3) amp

V1 = 12 (_cell voltage_) + (_voltage droped_) 1*(1/3)

V1 = 12 + 1/3 = (37/3) volt

V2 = 5(ohm)*i

V2 = 5*(1/3) = 5/3 volt

Similarly V3 = ( 2/3) volt

total resistance = 12 ohms